# Faster and Simpler Online Computation of String Net Frequency

### v2

@@ Abstract
An occurrence of a repeated substring $u$ in a string $S$ is called a net occurrence if extending the occurrence to the left or to the right decreases the number of occurrences to 1. The net frequency (NF) of a repeated substring $u$ in a string $S$ is the number of net occurrences of $u$ in $S$. Very recently, Guo et al. [SPIRE 2024] proposed an online $O(n\log\sigma)$-time algorithm that maintains a data structure of $O(n)$ space which answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time and reports all answers of the All-NF problem in $O(n\sigma^{2})$ time. Here, $n$ is the length of the input string $S$, $m$ is the query pattern length, and $\sigma$ is the alphabet size. The $\sigma^{2}$ term is a major drawback of their method since computing string net frequencies is originally motivated for Chinese language text processing where $\sigma$ can be thousands large. This paper presents an improved online $O(n\log\sigma)$-time algorithm, which answers Single-NF queries in $O(m\log\sigma)$ time and reports all answers to the All-NF problem in output-optimal $O(|\mathsf{NF^{+}}(S)|)$ time, where $\mathsf{NF^{+}}(S)$ is the set of substrings of $S$ paired with their positive NF values. We note that $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. In contract to Guo et al.’s algorithm that is based on Ukkonen’s suffix tree construction, our algorithm is based on Weiner’s suffix tree construction.
@@ 1 Introduction
[…]
@@ 2.2 String Net Frequency and Maximal Repeats
Since $\phi_{S}(w)\geq 1$, there exists a pair $(\alpha,\beta)$ of characters such that $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=1$. Since $\mathsf{occ}_{S}(w)\geq 2$, there is an occurrence $w=S[i..i+|w|-1]$ of $u$ such that $S[i-1]\neq\alpha$ or $i=1$. Similarly, there is an occurrence $w=S[j-|w|+1..j]$ of $w$ such that $S[j+1]\neq\beta$ or $j=|S|$. Thus $w\in\mathsf{LM}(S)\cap\mathsf{RM}(S)=\mathsf{M}(S)$. ∎
@@ 2.3 Online Computation of String Net Frequency
Increase the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)>\phi_{S}(w)\geq 1$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$.
@@
Add maximal repeats whose NFs become positive: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\notin\mathsf{NF^{+}}(S)$, $\phi_{S}(w)=0$, and $\phi_{aS}(w)\geq 1$, add $(w,\phi_{aS}(w))$ to the list $\mathcal{L}$.
@@
Decrease the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)<\phi_{S}(w)$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$ if $\phi_{aS}(w)\geq 1$, and remove $(w,\phi_{S}(w))$ from $\mathcal{L}$ if $\phi_{aS}(w)=0$.
@@
Consider string $S=\mathtt{ababbababcababbb\$}$ of which the set of maximal repeats is $\mathsf{M}(S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb}\}$. We have
  • $\Phi_{S}(\mathtt{bb})=\{(\underline{\mathtt{b}},\underline{\mathtt{\$}})\}$ and thus $\phi_{S}(\mathtt{bb})=1$, because $\mathsf{occ}_{S}(\mathtt{bb})=3$ and $\mathsf{occ}_{S}(\mathtt{\underline{b}bb})=\mathsf{occ}_{S}(\mathtt{bb\underline{\$}})=\mathsf{occ}_{S}(\mathtt{\underline{b}bb\underline{\$}})=1$;
@@
Consider string $\mathtt{b}S=\mathtt{bababbababcababbb\$}$ of which the set of maximal repeats is $\mathsf{M}(\mathtt{b}S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb},\mathtt{babab}\}$. For the new maximal repeat $\mathtt{babab}\in\mathsf{M}(\mathtt{b}S)\setminus\mathsf{M}(S)$ we have
@@
  • $\Phi_{\mathtt{b}S}(\mathtt{ababb})=\{(\mathtt{\underline{b}},\mathtt{\underline{a}}),(\mathtt{b},\mathtt{c})\}$ and thus $\phi_{\mathtt{b}S}(\mathtt{abab})=2$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb})=2$, $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb\underline{a}})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb\underline{a}})=1$, and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababbb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababbb})=1$. This is in Case (i);
@@ 2.4 Suffix Trees and Weiner Links
For convenience, we assume that each string $S$ terminates with a unique character \$currency-dollar\\$\$ that does not appear elsewhere in $S$.
@@
For any string $S$ of length $n$ that terminates with \$currency-dollar\\$\$, $\mathsf{STree}(S)$ has exactly $n$ leaves, and has at most $n-1$ internal nodes (including the root) since every non-leaf node is branching. By representing each edge label $x$ with an ordered pair $\langle i,j\rangle$ of integers such that $x=S[i..j]$, $\mathsf{STree}(S)$ can be stored in $O(n)$ space.
We define the suffix link of each non-root node $av$ of $\mathsf{STree}(S)$ with $a\in\Sigma$ and $v\in\Sigma^{*}$, by $\mathsf{slink}(av)=v$. For each node $v$ and $a\in\Sigma$, we also define the reversed suffix link (a.k.a. Weiner link) by ${\mathsf{W\_link}}_{a}({v})=avy$, where $y\in\Sigma^{*}$ is the shortest string such that $avy$ is a node of $\mathsf{STree}(S)$. ${\mathsf{W\_link}}_{a}({v})$ is undefined if $av\notin\mathsf{Substr}(S)$. A Weiner link ${\mathsf{W\_link}}_{a}({v})=avy$ is said to be hard if $y=\varepsilon$, and soft if $y\in\Sigma^{+}$.
@@
For any string $S$ of length $n>2$, the total number of hard and soft Weiner links in $\mathsf{STree}(S)$ is at most $43n-$ 4.
Throughout the description of our algorithm in Section 3, we will assume that the input string $S$ terminates with a unique end-marker \$currency-dollar\\$\$. We note that this assumption only gives little impact in our NF computation. By our key lemma (Lemma 3) from the next section, we have that there can exist at most one character $\alpha$ such that $(\alpha,\$)\in\Phi_{S}(w)$ for a repeat $w$ in $S$. Moreover, there can exist at most one repeat $w$ that can have \$currency-dollar\\$\$ in the second element of the pair $(\alpha,\$)\in\Phi_{S}(w)$: Let $w$ be the longest repeated suffix of $S[1..n-1]$ where $n=|S|$, and let $\alpha w$ be the shortest non-repeated suffix of $S[1..n-1]$ where $\alpha\in\Sigma$. Since $S[n]=\$$, we have $\mathsf{occ}_{S[1..n-1]}(w)=\mathsf{occ}_{S}(w)>1$ and $\mathsf{occ}_{S[1..n-1]}(\alpha w)=\mathsf{occ}_{S}(\alpha w)=1$. Since $\mathsf{occ}_{S}(w\$)=\mathsf{occ}_{S}(\alpha w\$)=1$, $(\alpha,\$)\in\Phi_{S}(w)$ holds. Note that no other suffixes of $S[1..n-1]$ can satisfy these properties.
@@ 3.1 Key Lemma
Since the total number of right-extensions of the maximal repeats is most $22n-$ 2 for any string of length $n\geq 2$ (c.f. [3]), the next lemma immediately follows from Lemma 1 and Corollary 1.
@@ 3.2 Recalling Weiner’s Suffix Tree Construction
Suppose that all the Weiner links have also been computed for the nodes of $\mathsf{STree}(S)$. Since our input string $S$ terminates with \$currency-dollar\\$\$ at its right end, all the suffixes of $S$ are represented by the leaves of $\mathsf{STree}(S)$.
@@
Figure 1: A snapshot of Weiner’s suffix tree construction algorithm. Solid arcs represent edges, and broken thick (resp. thin) arcs represent hard (resp. soft) Weiner links. Left: By climbing up the path from leaf $S$, we find the lowest ancestor $v$ that has a soft Weiner link ${\mathsf{W\_link}}_{a}({v})=ax$ that is labeled by $a$, and the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$ that is labeled by $a$. Right: Split the edge from node $az$ to $ax$ at the locus for $av$, and insert a new leaf for $aS$ from $av$. Then, the NF of node $u$ is increased by 1, since $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$.
In Weiner’s algorithm, we climb up the path from the leaf for $S$ and find the first ancestor $v$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined. See also Figure 1 for illustration. There are two cases:
@@ 3.3 Computing $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$
Suppose that we have built $\mathsf{STree}(S)$ for string $S$ of length $n$ augmented with soft and hard Weiner links, and that we have computed the list $\mathsf{NF^{+}}(S)$. For each node $u$ of $\mathsf{STree}(S)$ with $\phi_{S}(u)\geq 1$, each element $(u,\phi_{S}(u))\in\mathsf{NF^{+}}(S)$ is stored in $u$. By Lemma 1 (or alternatively by Lemma 4), we can store $\mathsf{NF^{+}}(S)$ for all nodes $u$ of $\mathsf{STree}(S)$ in total $O(n)$ space.
@@
Suppose $(\alpha,\beta)\in\Phi_{S}(w)$ holds. Then, by definition, $w$ is a repeat in $S$. Further, by Lemma 1 and by our assumption that $S$ terminates with a unique end-marker \$currency-dollar\\$\$, (1) $w$ is an internal branching node in $\mathsf{STree}(S)$. By assumption we have $\mathsf{occ}_{S}(w\beta)=1$, and thus (3) $w$ has an out-edge that begins with $\beta$ and leads to a leaf. Also, by assumption we have $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(\alpha w\beta)=1$. This implies that $\alpha w$ is on an edge leading to a leaf $\alpha wt$ with $t\in\Sigma^{+}$ and $t[1]=\beta$. Also, $\alpha w\beta$ is on the same edge if $|t|>1$, or it is a leaf if $|t|=1$. Since $w$ is preceded by $\alpha$ in $S$, $w$ has a Weiner link ${\mathsf{W\_link}}_{\alpha}({w})$ labeled $\alpha$. Since $\alpha w$ is on an edge, ${\mathsf{W\_link}}_{\alpha}({w})=\alpha wt$ with $|t|\geq 1$, and thus it is a soft Weiner link. This leads to (2).
@@
Existing maximal repeat $w$ in $aS$ that obtains a new left-extension $\alpha$ and an existing right-extension $\beta$ such that $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w)=\mathsf{occ}_{aS}(w\beta)=1$. Notice that $w\in\mathsf{M}(S)$ since $\phi_{S}(w)\geq 1$ for Case (i) in Section 2.3. We will later show that the parent $u$ of the leaf for $S$ is the only candidate for this type.
@@
New maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(w\beta)=1$. We will later show that $w\notin\mathsf{M}(S)$ and $w\in\mathsf{M}(aS)$ for any Type (ii) node $w$, and that the parent $av$ of the new leaf for $aS$ is the only candidate for this type.
@@
Existing maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(\alpha w)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-1, since $\phi_{S}(w)\geq 1$ for Case (iii) in Section 2.3. We will later show that the lowest ancestor $v$ of the leaf for $S$ that has a soft Weiner link labeled $a$ in $\mathsf{STree}(S)$ is the only candidate for this type.
@@
Existing maximal repeat $w$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(w\beta)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-2, since $\phi_{S}(w)=1$ for Case (iii) in Section 2.3. We will later show that the node $az$ that is the destination node of the hard Weiner link from $z$ labeled $a$ is the only candidate for this type.

### v3

@@ Abstract
An occurrence of a repeated substring $u$ in a string $S$ is called a net occurrence if extending the occurrence to the left or to the right decreases the number of occurrences to 1. The net frequency (NF) of a repeated substring $u$ in a string $S$ is the number of net occurrences of $u$ in $S$. Very recently, Guo et al. [SPIRE 2024] proposed an online $O(n\log\sigma)$-time algorithm that maintains a data structure of $O(n)$ space which answers Single-NF queries in $O(m\log\sigma+\sigma^{2})$ time and reports all answers of the All-NF problem in $O(n\sigma^{2})$ time. Here, $n$ is the length of the input string $S$, $m$ is the query pattern length, and $\sigma$ is the alphabet size. The $\sigma^{2}$ term is a major drawback of their method since computing string net frequencies is originally motivated for Chinese language text processing where $\sigma$ can be thousands large. This paper presents an improved online $O(n\log\sigma)$-time algorithm, which answers Single-NF queries in $O(m\log\sigma)$ time and reports all answers to the All-NF problem in output-optimal $O(|\mathsf{NF^{+}}(S)|)$ time, where $\mathsf{NF^{+}}(S)$ is the set of substrings of $S$ paired with their positive NF values. We note that $|\mathsf{NF^{+}}(S)|=O(n)$ always holds. In contrast to Guo et al.’s algorithm that is based on Ukkonen’s suffix tree construction, our algorithm is based on Weiner’s suffix tree construction.
@@ 1 Introduction
A repeat in string $S$ is a substring that occurs at least twice in $S$. An occurrence $[i..j]$ of a repeat $u$ in $S$ is said to be a net occurrence of $u$ if neither of its left-extension nor right-extension is a repeat, i.e., each of $au$, ubub, and aubaub occurs exactly once in $S$ where a=T[i1]a=T[i-1], $b=T[j+1]$, and $1<i\leq j<n$. The number of net occurrences of a repeat $u$ in a string $S$ is said to be the net frequency (NF) of $u$ in $S$. Computing the string net frequencies has applications in Chinese language text processing [11, 12, 15].
Efficient computation of string net frequencies has gained recent attention. Guo, Eades Wirth, and Zobel [8] proposed the first efficient algorithm for computing the string net frequencies (the GEWZ algorithm). The GEWZ algorithm works in an offline manner, namely it first reads the whole input string $S$, and builds a data structure of $O(n)$ space based on the suffix array [13] and the Burrows-Wheeler transform [5]. The data structure of the GEWZ algorithm takes $O(n)$ total space, and can be built in $O(n)$ time for integer alphabets of polynomial size in $n$, or in $O(n\log\sigma)$ time for general ordered alphabets of size $\sigma$. Then, the GEWZ algorithm answers Single-NF queries of reporting the net frequency of a given pattern in $S$ in $O(m+\log n+\sigma)$ time each [^1] Theorem 15 of the GEWZ algorithm [8] assumes that a pair (i,i+m1)(i,i+m-1) of positions for an occurrence [i..i+m1][i..i+m-1] of a pattern in text $S$ is given as a query, where $m$ is the pattern length. When a pattern is given as a string, then since their data structure is based on the suffix array, it requires $O(m+\log n)$ time for finding an occurrence.. They also considered the All-NF problem of reporting the net frequencies of all repeats in $S$ that have positive net frequencies, and showed how their data structure can answer the All-NF problem in $O(n)$ time. Ohlebusch, Büchler, and Olbrich [15] proposed another offline algorithm (the OBO algorithm) for the All-NF problem that works in $O(n)$ total time for integer alphabets of size $\sigma=O(n)$.
@@ 2.2 String Net Frequency and Maximal Repeats
Since $\phi_{S}(w)\geq 1$, there exists a pair $(\alpha,\beta)$ of characters such that $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=1$. Since $\mathsf{occ}_{S}(w)\geq 2$, there is an occurrence w=S[i..i+|w|1]w=S[i..i+|w|-1] of $w$ such that $S[i-1]\neq\alpha$ or $i=1$. Similarly, there is an occurrence w=S[j|w|+1..j]w=S[j-|w|+1..j] of $w$ such that $S[j+1]\neq\beta$ or $j=|S|$. Thus $w\in\mathsf{LM}(S)\cap\mathsf{RM}(S)=\mathsf{M}(S)$. ∎
@@ 2.3 Online Computation of String Net Frequency
Increase the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)>\phi_{S}(w)\geq 1$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$. Also, for each $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ such that $(w,\phi_{S}(w))\notin\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)>\phi_{S}(w)=0$, add $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$.
@@
Add new maximal repeats whose NFs become positive: For each $w\in\mathsf{M}(aS)\setminus\mathsf{M}(S)$ such that $(w,\phi_{S}(w))\notin\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)>\phi_{S}(w)=0$, add $(w,\phi_{aS}(w))$ to the list $\mathcal{L}$.
@@
Decrease the NFs of some existing maximal repeats: For each $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ such that $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ and $\phi_{aS}(w)<\phi_{S}(w)$, replace $(w,\phi_{S}(w))$ with $(w,\phi_{aS}(w))$ in the list $\mathcal{L}$ if $\phi_{aS}(w)\geq 1$, and remove $(w,\phi_{S}(w))$ from $\mathcal{L}$ if $\phi_{aS}(w)=0$.
@@
Consider string $S=\mathtt{ababbababcababbb\mathdollar}$ of which the set of maximal repeats is $\mathsf{M}(S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb}\}$. We have
[…]
@@
Consider string $\mathtt{b}S=\mathtt{bababbababcababbb\mathdollar}$ of which the set of maximal repeats is $\mathsf{M}(\mathtt{b}S)=\{\mathtt{b},\mathtt{ab},\mathtt{bb},\mathtt{bab},\mathtt{abab},\mathtt{ababb},\mathtt{babab}\}$. For the new maximal repeat $\mathtt{babab}\in\mathsf{M}(\mathtt{b}S)\setminus\mathsf{M}(S)$ we have
@@
  • $\Phi_{\mathtt{b}S}(\mathtt{ababb})=\{(\mathtt{\underline{b}},\mathtt{\underline{a}}),(\mathtt{c},\mathtt{b})\}$ and thus $\phi_{\mathtt{b}S}(\mathtt{abab})=2$, because $\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb})=2$, $\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababb\underline{a}})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{\underline{b}ababb\underline{a}})=1$, and $\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{ababbb})=\mathsf{occ}_{\mathtt{b}S}(\mathtt{cababbb})=1$. This is in Case (i);
@@ 2.4 Suffix Trees and Weiner Links
[…]
@@
[…]
We define the suffix link of each non-root node $av$ of $\mathsf{STree}(S)$ with $a\in\Sigma$ and $v\in\Sigma^{*}$, by $\mathsf{slink}(av)=v$. For each node $v$ and $a\in\Sigma$, we also define the reversed suffix link (a.k.a. Weiner link) by ${\mathsf{W\_link}}_{a}({v})=avy$, where $y\in\Sigma^{*}$ is the shortest string such that avyavy is a node of $\mathsf{STree}(S)$. ${\mathsf{W\_link}}_{a}({v})$ is undefined if $av\notin\mathsf{Substr}(S)$. A Weiner link ${\mathsf{W\_link}}_{a}({v})=avy$ is said to be hard if $y=\varepsilon$, and soft if $y\in\Sigma^{+}$.
@@
For any string $S$ of length $n>2$, the total number of hard and soft Weiner links in $\mathsf{STree}(S)$ is at most 3n43n-4.
[…]
@@ 3.1 Key Lemma
Since the total number of right-extensions of the maximal repeats is most 2n22n-2 for any string of length $n\geq 2$ (c.f. [3]), the next lemma immediately follows from Lemma 1 and Corollary 1.
@@ 3.2 Recalling Weiner’s Suffix Tree Construction
[…]
@@
Figure 1: A snapshot of Weiner’s suffix tree construction algorithm. Solid arcs represent edges, and broken thick (resp. thin) arcs represent hard (resp. soft) Weiner links. Left: By climbing up the path from leaf $S$, we find the lowest ancestor $v$ that has a soft Weiner link ${\mathsf{W\_link}}_{a}({v})=ax$ that is labeled by $a$, and the lowest ancestor $z$ that has a hard Weiner link ${\mathsf{W\_link}}_{a}({z})=az$ that is labeled by $a$. Right: Split the edge from node $az$ to axax at the locus for $av$, and insert a new leaf for aSaS from $av$. If the parent $u$ of $S$ is not $v$, then the NF of the parent $u$ of $S$ is increased by 1, since $(a,b)\in\Phi_{aS}(u)\setminus\Phi_{S}(u)$.
In Weiner’s algorithm, we climb up the path from the leaf for $S$ and find the lowest ancestor $v$ such that the Weiner link ${\mathsf{W\_link}}_{a}({v})$ with the new character $a$ is defined. See also Figure 1 for illustration. There are two cases:
@@ 3.3 Computing $\mathsf{NF^{+}}(aS)\triangle\mathsf{NF^{+}}(S)$
Suppose that we have built $\mathsf{STree}(S)$ for string $S$ of length $n$ augmented with soft and hard Weiner links, and that we have computed the list $\mathsf{NF^{+}}(S)$. For each node $w$ of $\mathsf{STree}(S)$ with $\phi_{S}(w)\geq 1$, $(w,\phi_{S}(w))\in\mathsf{NF^{+}}(S)$ is stored in $w$. By Lemma 1 (or alternatively by Lemma 4), we can store $\mathsf{NF^{+}}(S)$ for all nodes $w$ of $\mathsf{STree}(S)$ in total $O(n)$ space.
@@
[…]
@@
Existing maximal repeat $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ in $aS$ that obtains a new left-extension $\alpha$ and an existing right-extension $\beta$ such that $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w)=\mathsf{occ}_{aS}(w\beta)=1$. We will later show that the parent $u$ of the leaf for $S$ is the only candidate for this type.
@@
New maximal repeat $w\in\mathsf{M}(aS)\setminus\mathsf{M}(S)$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(\alpha w\beta)=\mathsf{occ}_{aS}(w\beta)=1$. We will later show that the parent $av$ of the new leaf for $aS$ is the only candidate for this type.
@@
Existing maximal repeat $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(\alpha w)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-1, since $\phi_{S}(w)\geq 1$ for Case (iii) in Section 2.3. We will later show that the lowest ancestor $v$ of the leaf for $S$ that has a soft Weiner link labeled $a$ in $\mathsf{STree}(S)$ is the only candidate for this type.
@@
Existing maximal repeat $w\in\mathsf{M}(aS)\cap\mathsf{M}(S)$ in $aS$ that has a left-extension $\alpha$ and a right-extension $\beta$ with $\mathsf{occ}_{S}(\alpha w)=\mathsf{occ}_{S}(w\beta)=\mathsf{occ}_{S}(\alpha w\beta)=1$ and $\mathsf{occ}_{aS}(w\beta)>1$. Notice that $w\in\mathsf{M}(S)$ holds for each node $w$ of Type (iii)-2, since $\phi_{S}(w)\geq 1$ for Case (iii) in Section 2.3. We will later show that the node $az$ that is the destination node of the hard Weiner link from $z$ labeled $a$ is the only candidate for this type.